OpenStudy (anonymous):

simplify this expression: 3 times (b-(-7)

OpenStudy (zzr0ck3r):

hint: \(3(a-(-b))=3(a+b)\)

OpenStudy (zzr0ck3r):

ie. b-(-7)= b+7